# Active Alignments of Lens Systems with Reinforcement Learning*

### v1

@@ I Introduction
Figure 1: Schematic presentation of a single alignment step, where a lens system consisting of three single lenses (left) has to be positioned relative to an optical sensor (right).
@@
  • We formulate active lens alignment problems as POMDP tangible by RL algorithms.
  • We introduce the Python framework relign [^1] Code under https://github.com/hs-kempten/relign. for simulating active alignment scenarios using the physically based rendering framework Mitsuba [21].
  • We provide an interface compatible with the Gymnasium API [22] for benchmarking state-of-the-art methods on representative alignment tasks, including RL models and black-box optimization techniques.
@@
Our approach not only speeds up the optical alignment process significantly, it also renders the need to design hand-crafted features obsolete.
@@ II-A Problem Formulation
An optical alignment process joins multiple components to maximize optical performance. In this work, the component to be aligned is a lens system consisting of a fixed number of single lenses $L=(L_{1},\ldots,L_{k})$ that need to be positioned relative to an optical sensor. This situation is typical when manufacturing cameras. The goal of an alignment is to move $L$ to a position $s=(x,y,z,R_{x},R_{y},R_{z})\in\mathbb{R}^{6}$ relative to a sensor, typically with an automated robotic alignment system, such that the optical performance is maximal.
The main challenge is that in each alignment, the process has to adjust to new conditions, mainly due to variances within $L$ and when gripping $L$. The first type of variances is a randomized offset arising from the placement of $L$ in the alignment station. This can be modeled by a randomized starting position $W_{\mathrm{off}}\in\mathbb{R}^{6}$ representing a random translation and rotation offset. When connecting to the robotic alignment system, the movement is typically not optimal, meaning that when a movement $a\in\mathbb{R}^{6}$ in a state $s$ is executed, the new state is not $s+a$ but the slightly distorted state $s+W_{\mathrm{dist}}\cdot a$ with $W_{\mathrm{dist}}\in\mathbb{R}^{6\times 6}$. Moreover, the new state can be clipped, for instance because a boundary condition is met. We refer to Section Section IV-B for more details on how $W_{\mathrm{dist}}$ is constructured and how the boundary behavior is modeled. Another type of variances comes from the production of $L$ itself, which cannot be changed by the process during the alignment. That is, each single lens $L_{i}$ in $L$ has an individual tilt and position offset in comparison to the ideal lens system (see Figure 2). We denote the offsets of each lens within $L$ by $W_{\mathrm{L}}\in\mathbb{R}^{k\times 6}$. Other variances not considered in this work are dispersions in the geometries of the single lenses, like their curvature. Here, the variances $W=(W_{\mathrm{off}},W_{\mathrm{dist}},W_{\mathrm{L}})$ characterize an alignment completely and we assume that these are sampled from an unknown distribution $W\sim\rho$. We further assume that the variances $W$ are latent to the alignment process, i.e., cannot be measured while aligning. As a result, it is not possible to directly compute the position where optical performance is optimal using the physical equations of e.g. geometrical optics.
Refer to caption Refer to caption $W_{\mathrm{L}}=0$$W_{\mathrm{L}}\neq 0$
Figure 2: Impact of $W_{\mathrm{L}}$ on lenses. Here, the arrows depict normal vectors of the lenses to visualize noise in the rotations.
@@ II-B Performance Measurements
To quantify the optical performance at a given state $s$, collimated light is sent through $L$ and measured at the sensor with width $w$ and height $h$, yielding a high-dimensional image $O_{W}(s)\in\mathbb{R}^{w\times h}$ of light intensities. The sensor output $O_{W}(s)$ is noisy, and retrieving $O_{W}(s)$ multiple times for the same position $s$ always yields slightly different observations. We call this noise sensor variance. In many industrial applications, hand-crafted scalar features are extracted from $O_{W}(s)$, typically involving the optical transfer functions, for which quality bounds have to be reached during the alignment. Here, however, we study problems where a generic reference output $O^{*}$ independent of $W$ is given such that the alignment task is the identification of a state $s^{*}$ where the sensor output matches the reference, i.e., $O_{W}(s^{*})\approx O^{*}$. The reference pattern can be considered as $O^{*}=\mathbb{E}[O_{W^{*}}(s^{*})]$, that is the sensor output without noise $W^{*}=0$ in the positioning of $L$ and its lenses at the optimal position $s^{*}$. In practice, the choice of the light field used for creating a test pattern at sensor level is dictated by the specific application requirements, such as whether the camera needs to achieve sharp focus in the near or far field. A common example is the Siemens star [23], which is widely used for evaluating optical performance of digital cameras [24]. Numerous methods exist to measure the quality of an optical image based on projected patterns such as a pattern of Siemens stars and slanted edges as described in ISO 12233 [24]. Given a reference pattern, the optimization problem
$\operatorname*{arg\,min}_{s\in S}\|O_{W}(s)-O^{*}\|$ (1)
[…]
Figure 3: Visualization of $(s_{i},s_{j})\to\min_{s_{k},k\neq i,k\neq j}\|O_{W}(s)-O^{*}\|$ for each tuple $\{i,j\}\subset\{x,y,z,R_{x},R_{y}\}$.
@@ II-C Alignment Algorithms
As the variances $W$ are unobservable by the alignment process, most active alignment algorithms explore the state space $S$ iteratively by selecting states and searching for directions where the score $\|O_{W}(s)-O^{*}\|$ attains its minimum. Once positioned at a new state, a new sensor signal is observed. An alignment algorithm computes a series of subsequent actions $a_{1},\ldots,a_{n}$, step by step, starting from a randomized initial position $s_{0}=W_{\mathrm{off}}$. Each action is an element of a set of allowed actions $A\subset\mathbb{R}^{6}$ and generates a sequence of states $s_{i}=s_{0}+W_{\mathrm{dist}}(a_{1}+\ldots+a_{i})$ such that $s_{n}$ is optimal, i.e., $\|O_{w}(s_{n})-O^{*}\|\leq\theta$. The computation of $a_{i}$ must be based on a subset of the $i-1$ many images $O_{W}(s_{1}),\ldots,O_{W}(s_{i-1})$ obtained so far. In Section V-B, we state the alignment algorithms used in our benchmark study.
III Approach
III-A Alignments as a POMDP
In this section, we describe how RL algorithms can be used as alignment algorithms as defined in Section II-C. Specifically, we consider active alignments as an episodic POMDP, where each episode is the alignment of a given lens system $L$. As defined in Section II, $S$ represents the set of states, and $A$ denotes the set of actions. In the alignment of $L$, including variances $W=(W_{\mathrm{off}},W_{\mathrm{dist}},W_{\mathrm{L}})$, selecting $a$ at $s$ results in the new state $s^{\prime}=s+W_{\mathrm{dist}}\cdot a$, yielding a reward $R(s,s^{\prime})$. A detailed view on different reward functions used is provided in Section III-B. The state $s\in S$ cannot be directly observed. Instead, only the high-dimensional sensor output $O_{W}(s)$ is given, which is controlled by a conditional probability density function depending on $s$. We explain in Section IV-A how an image $O_{W}(s)$ is sampled from the sensor at a given state $s$. For given $W$, an episode ends once a terminal state, that is an optimal state from $\{s\in S:\|O_{W}(s)-O^{*}\|\leq\theta\}$, or an upper limit of steps $T\in\mathbb{N}$ is reached. The goal is to find a policy $\pi$ that maps observations to actions in a way that maximizes the accumulated observed reward. More formally, given the observation $O_{W}(s)$ for $s$, the next state is $s+W_{\mathrm{dist}}\cdot\pi(O_{W}(s))$. Starting from an initial state $s_{0}$, this combined dynamics of sampled observations from the sensor and generated actions by the policy $\pi$ yields a trajectory $(s_{0},\ldots,s_{k})$ of subsequent states with $s_{i+1}=s_{i}+W_{\mathrm{dist}}\cdot\pi(O_{W}(s_{i}))$ where the last state $s_{k}$ is either optimal or $k=T$ holds. The goal is to find a policy such that
@@ III-A Alignments as a POMDP
III-B Rewards
To train effective agents for an alignment task, the reward function $R$ has to be designed in a way that the optimization task from Equation (2) is solved in a minimal number of steps. We consider two distinct reward functions: The first, $R_{\mathrm{step}}$, assigns a fixed penalty of $-1$ for each step taken, independent of the states and their distance to the goal, i.e., $R_{\mathrm{step}}(s,s^{\prime})=-1$. The second approach involves rewarding each step based on the extent to which the action brings the system closer to the optimal state. More formally, we define
$R_{\mathrm{pot}}(s,s^{\prime})=\|O_{W}(s)-O^{*}\|-\|O_{W}(s^{\prime})-O^{*}\|.$
That is, a larger reward is obtained if $s^{\prime}$ is closer to the optimal state than $s$. Note that $R_{\mathrm{pot}}$ is a potential-based reward [25], that means that if $\gamma=1$, the reward of a trajectory $(s_{0},\ldots,s_{k})$ satisfies $\sum_{i=0}^{k-1}R_{\mathrm{pot}}(s_{i},s_{i+1})=R_{\mathrm{pot}}(s_{0},s_{k})$.
@@ IV Simulating Alignments
One main challenge in simulating realistic optical alignments is to accurately model how a sensor measures light emitted from a source and propagated through optical lenses. Here, we use Mitsuba3, a physically based rendering engine for forward and inverse light-transport simulation. This not only allows calculating light intensities $O_{W}(s)$ measured at the sensor (see Section IV-A) for a concrete position $s$, but also changing $s$ dynamically (see Section IV-B). All Mitsuba scenes consist of a sensor, $k$ many identical biconvex single lenses, and a light source.
@@ IV-A Sensor Outputs
The Mitsuba scene emulates collimated light using a binary environment map with a centered rectangular light source. Starting at the sensor, Mitsuba traces light rays backwards that pass through the lenses. We use an irradiance meter as sensor, which measures the incident power per unit area over a predefined shape. In our setup, the sensor shape is $0.5\times 0.5$ in Mitsuba space coordinates. To solve the high-dimensional problem of rendering $O_{W}(s)$ numerically, Mitsuba employs Monte Carlo integration, drawing $s_{c}$ many samples from a uniform distribution. As a consequence, the sensor output $O_{W}(s)$ is different when rendering the same position $s$ with different seeds. This leads to the probability density function of the POMDP as described in Section III-A. An example of the resulting sensor output is shown in Figure 1. The measurements are interpreted as a grayscale image and stored without any post-processing.
@@ IV-B Position Changes
Without loss of generality, we assume that the set of all possible states is the unit interval $[0,1]^{6}$. We place $k$ many lenses in the lens system $L$ in a way that $s^{*}:=(0.5,\ldots,0.5)\in\mathbb{R}^{6}$ is the optimal position. Upon initialization, we sample $W_{\mathrm{L}}$ from a normal distribution and reposition the single lenses within $L$ accordingly. Afterward, a starting vector $W_{\mathrm{off}}\in[0,1]^{6}$ is sampled uniformly which defines the initial state $s_{0}:=W_{\mathrm{off}}$. The movement distortion matrix is constructed as $W_{\mathrm{dist}}=I_{6}+{\epsilon}\in\mathbb{R}^{6\times 6}$ where each coordinate in ${\epsilon}\in\mathbb{R}^{6\times 6}$ is sampled from a normal distribution. The positioning of the lens system can be varied by setting an action $a\in\mathbb{R}^{6}$ leading to an update $s^{\prime}=s+W_{\mathrm{dist}}\cdot a$. To ensure that $s^{\prime}$ stays within $[0,1]^{6}$, each coordinate of $s+W_{\mathrm{dist}}\cdot a$ is clipped into $[0,1]$ before updating the Mitsuba scene. Subsequently, the sensor output $O_{W}(s^{\prime})$ is generated as described in Section IV-A.

### v2

@@ I Introduction
Figure 1: Schematic presentation of a single alignment step, where a lens system consisting of four single lenses (left) has to be positioned relative to an optical sensor (right).
@@
  • We formulate general active alignment problems as POMDP tangible by RL algorithms. To the best of our knowledge, this is the first work in doing so.
  • We introduce the modular Python framework relign [^1] Code under https://github.com/hs-kempten/relign. for simulating active alignment scenarios using the physically based rendering framework Mitsuba [21]. Its interface is compatible with the Gymnasium API [22] and allows effective benchmarking of state-of-the-art methods on representative alignment tasks.
@@
Finding accurate and comparable benchmarks for active alignments is hard, not only due to privacy constraints of companies maintaining them as part of their core know-how, but also due to the fact that optical products are very diverse and have different requirements. Here, our framework provides a first step towards a common benchmarking environment for active alignments. Moreover, we argue that using RL for alignments not only has the potential to significantly speed up the optical alignment process but also renders the need for designing hand-crafted features obsolete.
@@ II Active Alignments of Lens Systems
Automated robotic alignment systems often rely on high-precision motion platforms such as hexapods or piezo stages, which provide movement in all six degrees of freedom. These mechanisms enable sub-micrometer adjustments, allowing the robot to finely correct tilt, shift, and focus errors until the desired alignment is reached. In this section, we formalize the active alignment problem and introduce the main challenges faced during the alignment process.
@@ II-A Problem Formulation
An optical alignment process joins multiple components to maximize optical performance. In this work, the component to be aligned is a lens system consisting of a fixed number of single lenses $L=(L_{1},L_{2},L_{3},L_{4})$ that need to be positioned relative to an optical sensor. This situation is typical when manufacturing cameras. The goal of an alignment is to move $L$ to a position $s=(x,y,z,R_{x},R_{y},R_{z})\in\mathbb{R}^{6}$ relative to a sensor, typically with an automated robotic alignment system, such that the optical performance is maximal.
The main challenge is that in each alignment, the process has to adjust to new conditions, mainly due to variances within $L$ and when gripping $L$. The first type of variance is a randomized offset arising from the initial placement of $L$ in the alignment station. This can be modeled by a randomized starting position $W_{\mathrm{off}}\in\mathbb{R}^{6}$ representing a random translation and rotation offset. When connecting to the robotic alignment system, the movement is typically not optimal, meaning that when a movement $a\in\mathbb{R}^{6}$ in a state $s$ is executed, the new state is not $s+a$ but the slightly distorted state $s+W_{\mathrm{dist}}\cdot a$ with $W_{\mathrm{dist}}\in\mathbb{R}^{6\times 6}$. Moreover, the new state can be clipped, for instance because a boundary condition is met. We refer to Section IV-B for more details on how $W_{\mathrm{dist}}$ is constructed and how the boundary behavior is modeled. Another type of variance comes from the production of $L$ itself, which cannot be changed by the process during the alignment. That is, each single lens $L_{i}$ in $L$ has an individual tilt and position offset in comparison to the ideal lens system. We denote the offsets of each lens within $L$ by $W_{\mathrm{L}}\in\mathbb{R}^{k\times 6}$. Other variances not considered in this work are dispersions in the geometries of the single lenses, like their curvature. Here, the variances $W=(W_{\mathrm{off}},W_{\mathrm{dist}},W_{\mathrm{L}})$ characterize an alignment completely and we assume that these are sampled from an unknown distribution $W\sim\rho$. We further assume that the variances $W$ are latent to the alignment process, i.e., cannot be measured while aligning. As a result, it is not possible to directly compute the position where optical performance is optimal using the physical equations of e.g. geometrical optics.
@@ II-B Performance Measurements
To quantify the optical performance at a given state $s$, collimated light is sent through $L$ and measured at the sensor with width $w$ and height $h$, yielding a high-dimensional image $O(s,W)\in\mathbb{R}^{w\times h}$ of light intensities. The sensor output $O(s,W)$ is noisy, and retrieving $O(s,W)$ multiple times for the same position $s$ always yields slightly different observations. We call this noise sensor variance. In many industrial applications, hand-crafted scalar features are extracted from $O(s,W)$, typically involving the optical transfer functions, for which quality bounds have to be reached during the alignment. Here, however, we study problems where a generic reference output $O^{*}$ independent of $W$ is given such that the alignment task is the identification of a state $s_{W}$ where the sensor output matches the reference, i.e., $O(s_{W},W)\approx O^{*}$. The reference pattern can be considered as $O^{*}=\mathbb{E}[O_{W}(s_{W})]$, that is the sensor output without noise $W^{*}=0$ in the positioning of $L$ and its lenses at the respective optimal position $s_{W}$. In practice, the choice of the light field used for creating a test pattern at sensor level is dictated by the specific application requirements, such as whether the camera needs to achieve sharp focus in the near or far field. A common example is the Siemens star [23], which is widely used for evaluating optical performance of digital cameras [24]. Numerous methods exist to measure the quality of an optical image based on projected patterns such as a pattern of Siemens stars and slanted edges as described in ISO 12233 [24]. Given a reference pattern, the optimization problem
$\operatorname*{arg\,min}_{s\in S}\|O(s,W)-O^{*}\|$ (1)
has to be solved for a given situation $W\sim\rho$, where $S\subset\mathbb{R}^{6}$ denotes the set of allowed positions. Here, $\|\cdot\|$ denotes a vector norm, like the Euclidean distance, and we interpret the input matrices as vectors. Figure 2 shows some 2D-projections of this optimization problem for the alignment situations introduced in Section VI. Typically, a threshold $\theta\in\mathbb{R}_{\geq 0}$ is given such that any $s$ that satisfies $\|O(s,W)-O^{*}\|\leq\theta$ is considered as optimal. However, as $s$ can only compensate positional offsets of $L$ given by $W_{\mathrm{off}}$ under distortions given by $W_{\mathrm{dist}}$ and not the variances $W_{\mathrm{L}}$ inside $L$, some lens systems may have $\|O(s,W)-O^{*}\|>\theta$ for all positions $s$.
Figure 2: Visualization of the two-dimensional projection $(s_{i},s_{j})\mapsto\|O(s^{*}-s_{i}e_{i}-s_{j}e_{j},W^{*})-O^{*}\|$ for each tuple $\{i,j\}\subset\{x,y,z,R_{x},R_{y}\}$, that is, all except two dimensions are fixed to the optimal values.
@@ II-C Alignment Algorithms
As the variances $W$ are unobservable by the alignment process, most active alignment algorithms explore the state space $S$ iteratively by selecting states and searching for directions where the score $\|O(s,W)-O^{*}\|$ attains its minimum. Once positioned at a new state, a new sensor signal is observed. An alignment algorithm computes a series of subsequent actions $a_{1},\ldots,a_{n}$, step by step, starting from a randomized initial position $s_{0}=W_{\mathrm{off}}$. Each action is an element of a set of allowed actions $A\subset\mathbb{R}^{6}$ and generates a sequence of states $s_{i}=s_{0}+W_{\mathrm{dist}}(a_{1}+\ldots+a_{i})$ such that $s_{n}$ is optimal, i.e., $\|O(s_{n},W)-O^{*}\|\leq\theta$. The computation of $a_{i}$ must be based on a subset of the i1i-1 many images $O(s_{1},W),\ldots,O(s_{i-1},W)$ obtained so far. In Section VI-A, we state the alignment algorithms used in our benchmark study.
III Alignments as a POMDP
In this section, we describe how RL algorithms can be used as alignment algorithms as defined in Section II-C. Specifically, we consider active alignments as an episodic POMDP, where each episode is the alignment of a given lens system $L$. Here, a state $(s,W)$ is decomposed in a position $s\in S$ and variances $W=(W_{\mathrm{off}},W_{\mathrm{dist}},W_{\mathrm{L}})$ as defined in Section II and the set $A$ denotes the set of actions. Selecting $a$ at $(s,W)$ results in the new state $(s^{\prime},W)$ with $s^{\prime}=s+W_{\mathrm{dist}}\cdot a$, yielding a reward $R(s,s^{\prime})$. As of the state representing the variances stays constant through an episode, this MDP is also an contextual MDP in the sense of [25, Section 3]. The state $(s,W)$ cannot be directly observed.
Instead, only the high-dimensional sensor output $O(s,W)$ is given, which is controlled by a conditional probability density function depending on $s$. We explain in Section IV-A how an image $O(s,W)$ is sampled from the sensor at a given state $s$. For given $W$, an episode ends once a terminal state, that is an optimal state from $\{s\in S:\|O(s,W)-O^{*}\|\leq\theta\}$, or an upper limit of steps $T\in\mathbb{N}$ is reached. The goal is to find a policy $\pi$ that maps observations to actions in a way that maximizes the accumulated observed reward. More formally, given the observation $O(s,W)$ for $s$, the next state is $s+W_{\mathrm{dist}}\cdot\pi(O(s,W))$. Starting from an initial state $s_{0}$, this combined dynamics of sampled observations from the sensor and generated actions by the policy $\pi$ yields a trajectory $(s_{0},\ldots,s_{k})$ of subsequent states with $s_{i+1}=s_{i}+W_{\mathrm{dist}}\cdot\pi(O(s_{i},W))$ where the last state $s_{k}$ is either optimal or $k=T$ holds. The goal is to find a policy such that
@@ III Alignments as a POMDP
To train effective agents for an alignment task, the reward function $R$ has to be designed in a way that the optimization task from Equation (2) is solved in a minimal number of steps. During training time of a optical system $L$ with variances $W$, agents can be provided with dense and insightful reward signals, for instance containing information of the distance to the desired optimum $s_{W}$, which need to be known beforehand. An ideal reward signal would be $\|s-s_{W}\|$ which, however, requires the knowledge of $s_{W}$. In our simulation and during train time, this could be computed, for instance using a conventional alignment algorithm, but this would slow down training time as this needs to be done for each training episode anew. Instead, we use an indirect signal which gives the distance to $s^{*}$, i.e. the optimal position if $W=0$:
$R_{\text{dist}}(s)=-\|s-s^{*}\|.$
However, $s_{W}$ may differ significantly from $s^{*}$, especially in degenerated situations where $W$ is large. Another reward signal could be to use the reference pattern $O^{*}$, which is known beforehand, like
$R_{\text{pattern}}(s,W)=-\|O(s,W)-O^{*}\|.$
Yet, even at the optimal position, $O(s_{W},W)$ may differ significantly from $O^{*}$ and we found that for complex patterns, the many local minima of $R_{\text{pattern}}$ make it hard for agents to learn good policies from that reward signal alone. Thus, we considered a combination of both rewards signals
$R(s,W)=\begin{cases}R_{\text{pattern}}(s,W)+R_{\text{dist}}(s),\quad&\text{if }\|s-s^{*}\|\leq C\\ R_{\text{pattern}}(s,W),&\text{otherwise}\end{cases}.$
This has the benefit that far away from $s^{*}$, and hence $s_{W}$, direction to the optimal region is enforced. Once being sufficiently close to $s^{*}$, the agents are purely guided by the pattern distance of the lens system at hand.
@@ IV Simulating Alignments
One main challenge in simulating realistic optical alignments is to accurately model how a sensor measures light emitted from a source and propagated through optical lenses. Here, we use Mitsuba3, a physically based rendering engine for forward and inverse light-transport simulation. This not only allows calculating light intensities $O(s,W)$ measured at the sensor (see Section IV-A) for a concrete position $s$, but also changing $s$ dynamically (see Section IV-B). All Mitsuba scenes consist of a sensor, a lens system with four single lenses, and a light source. As one of our contribution is the open-source framework relign, we put high engineering effort in providing a realistic simulation environment and keeping the sim-to-real gap small.
@@ IV-A Sensor Outputs
The Mitsuba scene emulates collimated light from a rectangular area emitter source and a specific reference pattern as binary bitmap. The light source is so far away that it can be considered infinity. Starting at the sensor, Mitsuba traces light rays backwards that pass through the lenses. We use an irradiance meter as sensor, which measures the incident power per unit area over a predefined shape. In our setup, the sensor shape is $2\times 2$ in Mitsuba space coordinates. To solve the high-dimensional problem of rendering $O(s,W)$ numerically, Mitsuba employs Monte Carlo integration, drawing $s_{c}$ many samples from a uniform distribution. As a consequence, the sensor output $O(s,W)$ is different when rendering the same position $s$ with different seeds. This leads to the probability density function of the POMDP as described in Section III. An example of the resulting sensor output is shown in Figure 1. The measurements are interpreted as a grayscale image and stored without any post-processing.
@@ IV-B Position Changes
Without loss of generality, we assume that the set of all possible states is the unit interval $0,1]^{6}$. We place our lens system $L$ in a way that $s^{*}:=(0.5,\ldots,0.5)\in\mathbb{R}^{6}$ is the optimal position. Upon initialization, we sample $W_{\mathrm{L}}$ from a normal distribution and reposition the single lenses within $L$ accordingly. Afterward, a starting vector $W_{\mathrm{off}}\in[0,1]^{6}$ is sampled uniformly which defines the initial state $s_{0}:=W_{\mathrm{off}}$. The movement distortion matrix is constructed as $W_{\mathrm{dist}}=I_{6}+{\epsilon}\in\mathbb{R}^{6\times 6}$ where each coordinate in ${\epsilon}\in\mathbb{R}^{6\times 6}$ is sampled from a normal distribution. The positioning of the lens system can be varied by setting an action $a\in\mathbb{R}^{6}$ leading to an update $s^{\prime}=s+W_{\mathrm{dist}}\cdot a$. To ensure that $s^{\prime}$ stays within $0,1]^{6}$, each coordinate of $s+W_{\mathrm{dist}}\cdot a$ is clipped into $[0,1]$ before updating the Mitsuba scene. Subsequently, the sensor output $O(s^{\prime},W)$ is generated as described in Section IV-A.